# Mapping the positions of Two-Level-Systems on the surface of a superconducting transmon qubit

### v1

@@ Introduction
The nature of two-level tunneling systems (TLS) in amorphous materials has been puzzling generations of physicists [1]. Today, TLS are recognized as the primary source of decoherence in superconducting qubits. A type of TLS that was well-studied in glasses is thought to originate in the tunneling of a single or a few atoms between two slightly different locations in the disordered material as illustrated in Fig. 1a [2]. In superconducting circuits, amorphous surface oxides on electrodes and those used for tunnel barriers of qubit junctions are thus a known host for TLS [3, 4, 5, 6, 7, 8]. In addition, microfabrication techniques were shown to spoil the crystallinity of the substrate and to leave residuals of glassy photoresist [9, 10]. There is a variety of other models of TLS formation, and it remains unknown which types of TLS are limiting qubit coherence [11]. To obtain insight into the elusive microscopic TLS structure, atomistic modeling has gained in importance and was e.g. used to characterize TLS formed by Hydrogen interstitials[12], by dangling surface atoms [13], and by tunneling atoms in Josephson junctions [14, 15].
@@
The qubit was fabricated from aluminum on a sapphire substrate, using optical lithography and dry etching for the ground plane and qubit island, and eBeam-patterned Dolan bridges to form the tunnel junctions and their leads in a 3-angle shadow evaporation process that avoids unwanted stray junctions[42]. Details on sample fabrication and circuit parameters are found in Supplementary Note A. The sample was cooled to a temperature of 25 - 30 mK and measured in a standard setup as detailed in Supplementary Note B.
The qubit showed energy relaxation times $T_{1}$ between 5 to 8 µs at operation frequencies between 5 and 5.5 GHz. Similarly fabricated qubits without gate electrodes achieved only slightly longer $T_{1}$-times between 10 and 20 µs [42]. However, the observed $T_{1}$-time falls within the estimate range of the radiative loss via the capacitive coupling to the four electrodes, which were placed in close vicinity to the qubit island to enhance the spatial resolution in TLS localization. This loss channel can be mitigated with an improved design of the qubit and on-chip electrodes [43]. Further details on loss are discussed in Supplementary Note C.
@@
Figure 2: Finding the location of a TLS. a TLS spectroscopy using the protocol in the inset, to reveal the resonance frequencies of TLS from minima (dark pixels) in the resulting qubit population $P(|1\rangle$. In each segment, the voltage on a different electrode $\alpha..\delta$ is increased by 1V. The highlighted trace shows a TLS as it is tuned through its symmetry point (arrow) according to Eq. (1) as illustrated in the lower inset. The TLS’ response strengths $\gamma_{i}$ to the different electrodes are obtained by fitting such traces to Eq. (1), and provide information about the TLS’ distance to the electrodes. b Difference between the measured TLS response strength ratio $\gamma_{i}/\gamma_{j}$ and the corresponding simulated E-field ratio (colorscale) for the TLS observed in a. Minima (dark pixels) indicate possible TLS positions. Each panel shows data from a different electrode pair as marked in the legends together with the measured tuning ratio. c The colorscale shows the difference sum $\sigma$ (Eq. 3) over all 6 unique combinations of electrode pairs. The minimum (white circle) marks the most probable TLS position.
To measure the TLS’ coupling strengths to the four DC-electrodes, their resonances are traced by TLS spectroscopy while the voltages on the DC-electrodes are swept. As an example, Figure 2a shows the resonance of a TLS that was tuned through the symmetry point of its hyperbola Eq. (1), where in each segment the voltage on the indicated gate electrode $\alpha...\delta$ was stepwise increased to a total of 1 V while other voltages were kept constant. The slope of the hyperbola in the different segments then depends on the tuning strength of the TLS by the corresponding electrode. It is obtained by fitting such traces to Eq. (1), where the factor for the induced asymmetry energy $2\textbf{p}\cdot\textbf{E}$ is replaced by $\sum_{i}\gamma_{i}V_{i}$. Here, $i\in\{\alpha,\beta,\gamma,\delta\}$ indicates the electrode that is biased by the voltage $V_{i}$, and $\gamma_{i}$ are the fitting tuning strengths that contain information on the distance of the TLS to the corresponding electrodes.
@@
To be able to detect a TLS in qubit $T_{1}$-swap-spectroscopy, its (resonant) coupling strength to the qubit $g=\left(\frac{\Delta}{\hbar\omega_{\mathrm{TLS}}}\right)\ \mathbf{p}\cdot\mathbf{E_{\mathrm{rms}}}$ must be large enough to result in a measurable decrease of the qubit’s energy relaxation rate. Thus, TLS can only be detected in circuit regions where the qubit’s AC-electric field strength $|E_{\mathrm{rms}}|$ exceeds $g_{\mathrm{min}}/p_{\parallel}$. Assuming $\Delta/\hbar\omega_{\mathrm{TLS}}\approx 1$ (most strongly coupling TLS near their symmetry point) and a field-parallel dipole moment of $p_{\parallel}\approx 1e$Å [3, 4, 44, 36], the required minimum coupling strength $g_{\mathrm{min}}$ can be estimated from the energy relaxation rate of the resonantly coupled qubit-TLS system [4] $\Gamma_{1}=2\,(g_{\mathrm{min}}/\hbar)^{2}/\Gamma+\Gamma_{1,Q}$, where $\Gamma_{1,Q}$ is the energy relaxation rate of the isolated qubit, and $\Gamma=\Gamma_{1,\mathrm{TLS}}/2+\Gamma_{2,\mathrm{TLS}}+\Gamma_{1,\mathrm{Q}}/2+\Gamma_{2,\mathrm{Q}}$ is the sum of TLS and qubit energy relaxation and dephasing rates. The assumption that TLS are detected if they reduce the qubit’s $T_{1}$ time by a factor of $\kappa$ translates into a minimum coupling strength of $g_{\mathrm{min}}=\hbar\sqrt{\kappa\cdot\Gamma_{\mathrm{1,Q}}\,\Gamma/2}$ and corresponding minimum AC-electric qubit field strength $E_{\mathrm{min}}=g_{\mathrm{min}}/p_{\parallel}$. A plot of these relations can be found in Supplementary Note D. For a qubit $T_{1}=7\$µs, TLS coherence times in the range of $T_{\mathrm{1,TLS}}\approx T_{\mathrm{2,TLS}}\approx 0.1-2\,$µs [4, 45, 6, 32, 46, 47], and $\kappa=5\%$, we find a $E_{\mathrm{min}}$ range of $\approx 0.3-3$ V/m.
@@
Figure 3a shows the simulated magnitude of the qubit’s AC-electric field $|E_{\mathrm{rms}}|$. A cross section through the edges of ground plane and qubit island is shown in Fig 3b and illustrates that in our qubit sample, TLS can only be detected within an $\approx 1-2$ µm distance from the edge of qubit electrodes where $|E_{\mathrm{rms}}|>E_{\mathrm{min}}$. Since $E_{\mathrm{min}}$ decreases with the square root of the qubit’s $T_{1}$ time, more coherent qubits are affected by TLS in a wider area which includes more weakly coupled ones as illustrated in Supplementary Note D.
@@
To show that the electric fields from different on-chip electrodes are indeed sufficiently parallel, such that the dipole moment in Eq. (2) can be canceled in order to simplify the solution for possible TLS positions, in Fig. 3c we plot the electric fields’ $\vec{x}$- and $\vec{z}$-components along the white line in Fig. 3a. Near the center of the gap between ground and qubit island, the E-fields point in different directions. However, within the short distance to the electrode edges where TLS are detected in our experiment, all E-fields are well aligned since their vector products with the qubit field $\vec{E_{q}}$ approach unity as plotted in Fig. 3d.
@@
Figure 4: TLS locations and properties. a Map of the individual positions of detected surface-TLS (yellow circles). Most TLS ($\approx$ 59%) were found at the Josephson junctions leads (see lower inset for a zoom on the DC-SQUID). Near the edges of capacitor island and ground plane, 27% and 14% of TLS were identified, respectively. b Histograms and cumulative distributions (insets) of the TLS’ electric dipole moment component $p_{\parallel}$ (upper panel) and TLS-qubit coupling strengths $g/h$ (lower panel) as estimated from the TLS’ positions, their coupling strengths to the electrodes, and the local magnitude of the qubit AC field $E_{\mathrm{rms}}$. c Percentage of TLS identified on the SQUID vs. ground plane plus island, plotted as a function of the electric field threshold $E_{\mathrm{min}}$ that accounts for TLS observability in swap spectroscopy. The solid lines are the ratios of $E_{\mathrm{rms}}^{2}$ integrated over the regions of the SQUID (red) and ground plus island (blue) where $E_{\mathrm{rms}}>E_{\mathrm{min}}$. The dashed lines are a fit of these ratios to the observed TLS distribution, obtained by assuming that the TLS density in the SQUID area is enhanced by a factor of 2. d Determined median TLS electric dipole moment $p_{\parallel}$ as a function of $E_{\mathrm{min}}$.
@@ Results
To generate a map of the individual TLS positions on the qubit circuit, we swept the voltages on each of the four electrodes in a range between -60V and +100V and performed TLS spectroscopy in a 150-MHz wide window as shown in Fig. 2a. In total, 55 TLS were observed in a single qubit sample whose tuning strength to all four electrodes could be characterized (see Supplementary Note E for further details).
@@
To check the mapping procedure’s reliability, the analysis is repeated for various values of $E_{\mathrm{min}}$. The resulting percentages of TLS found on the SQUID leads vs. those on the qubit island and ground plane are plotted in Fig. 4c. These remain at an approximate ratio of 60:40 within the most probable range of $E_{\mathrm{min}}\approx 0.3..3$ V/m as estimated above, which backs the robustness of the method. Above a value of $E_{\mathrm{min}}\approx 6$ V/m, the procedure fails since it is forced to place all TLS on the DC-SQUID.
Notably, the observed TLS distribution points towards an excess density of TLS near the DC-SQUID leads. Since the probability to detect a TLS scales with the square of the local qubit field, the expected ratio of the numbers of TLS observed near the SQUID vs. ground plane and qubit island can be estimated by comparing the integrals of $E_{\mathrm{rms}}^{2}$ over these regions. Hereby, we limit the integrals to where $E_{\mathrm{rms}}>E_{\mathrm{min}}$ to account for TLS detectability in swap spectroscopy, which sets this calculation apart from the common participation ratio analysis [5] that includes also weakly coupled TLS. The result is plotted with solid red and blue lines in Fig. 4c, and predicts that most TLS (>50%) would be observed on either the qubit island or ground plane in the whole range of reasonable values $E_{\mathrm{min}}<4$ V/m, in stark contrast to the experimental data.
@@ Discussion
We have demonstrated a method to determine the individual positions and electric dipole moments of TLS defects in a transmon qubit. The majority ($\approx 58\%$) of observed strongly interfering TLS in the qubit sample were found to reside on the qubit’s DC-SQUID. This confirms that the leads of tunnel junctions, due to their large electric energy participation, are a critical component which can dominate qubit loss [48, 49, 47], and advocates for wire-tapering techniques to dilute the electric field [50].
Our results additionally indicate that the TLS density near the junction leads is enhanced by a factor of $\approx 2$. This may be attributed to their different fabrication technique as compared to the qubit capacitor, which can promote TLS formation in various ways. For example, junction formation by shadow evaporation and electron-beam lithography is associated with larger amounts of resist residuals and enhanced roughness of junction lead interfaces [10, 51]. The junctions were fabricated with a lift-off process which reportedly leaves excess residues [9, 49]. Also, thinner films showed a larger density of grain boundaries associated with enhanced oxygen diffusion [18].
@@
Moreover, control over the local DC-electric field enables one to actively suppress decoherence by tuning dominating TLS defects out of the qubit resonance [38, 39, 40]. For this, multiple electrodes provide independent control over TLS at different locations which enhances the ability to decouple the qubit from the decohering bath.
@@ Acknowledgements
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
The authors declare no competing interests.
@@
A summary of the qubit circuit parameters is given by Table S1.
@@
Table S1: Measured and simulated fabrication parameters of the studied qubit sample. Capacitance values were extracted with Ansys Maxwell. Junction ($J$) and SQUID resistances are averages from 4-probe measurements.
@@
$Q_{l}=\left(\frac{C_{\mathrm{tot}}}{C_{c}}\right)^{2}\frac{Z_{q}}{\mathrm{Re}(Z_{\mathrm{eff}})},$ (4)
where $C_{\mathrm{tot}}=C_{q}+\sum_{i}C_{q,i}$ is the total capacitance, and $Z_{q}=\sqrt{L_{q}/C_{\mathrm{tot}}}\approx 290\,\Omega$ is the qubit impedance estimated from the inductance of its two parallel Josephson junctions $L_{q}=0.5\cdot L_{J}=0.5\cdot h/4\pi eI_{c}$ each having a critical current of $I_{c}\approx 23$ nA as shown in Table S1. Assuming a real part impedance of $\mathrm{Re}(Z_{\mathrm{eff}})\approx 50\,\Omega$ for the coplanar gate electrode including its feed line results in $Q_{l}\approx 0.6$ million. The resulting loaded quality factor for all four electrodes is $Q_{\mathrm{tot}}=(4/Q_{l})^{-1}$ and limits the qubit’s energy relaxation time to $T_{1}=Q_{\mathrm{tot}}/2\pi f_{q}\approx 5\,$µs at a qubit resonance frequency of $f_{q}=5$ GHz.
@@
Supplementary Note D TLS detectability limits
@@
$\Gamma_{1}=\frac{2\,(g/h)^{2}\,\Gamma}{(\Gamma/2\pi)^{2}+\delta^{2}}+\Gamma_{\mathrm{1,q}}.$ (5)
@@
Solving for the coupling strength $g$ where Eq. (5) predicts that the qubit’s energy relaxation time at resonance $\delta=0$ is decreased by a factor $\kappa$ then results in the minimum coupling strength $g_{\mathrm{min}}=\hbar\sqrt{\kappa\cdot\Gamma_{\mathrm{1,q}}\,\Gamma/2}$ and corresponding minimum AC-electric qubit field strength $E_{\mathrm{min}}=g_{\mathrm{min}}/p_{\parallel}$.
@@
Supplementary Note E Data acquisition and TLS properties
@@
Supplementary Note F Resolution in TLS mapping

### v2

# arXiv:2511.05365v2[quant-ph] 03 Mar 2026

@@ Introduction
The nature of two-level tunneling systems (TLS) in amorphous materials has been puzzling generations of physicists [12]. Today, TLS are recognized as the primary source of decoherence in superconducting qubits. A type of TLS that was well-studied in glasses is thought to originate in the tunneling of a single or a few atoms between two slightly different locations in the disordered material as illustrated in Fig. 1a [21]. In superconducting circuits, amorphous surface oxides on electrodes and those used for tunnel barriers of qubit junctions are thus a known host for TLS [34, 1, 52, 30, 4, 17]. In addition, microfabrication techniques were shown to spoil the crystallinity of the substrate and to leave residuals of glassy photoresist [45, 20, 24]. There is a variety of other models of TLS formation, and it remains unknown which types of TLS are limiting qubit coherence [39]. To obtain insight into the elusive microscopic TLS structure, atomistic modeling has gained in importance and was e.g. used to characterize TLS formed by Hydrogen interstitials[26], by dangling surface atoms [53], and by tunneling atoms in Josephson junctions [19, 44].
@@ Methods
The qubit was fabricated from aluminum on a sapphire substrate, using optical lithography and dry etching for the ground plane and qubit island, and eBeam-patterned Dolan bridges to form the tunnel junctions and their leads (shown in Fig. 1d) in a 3-angle shadow evaporation process that avoids unwanted stray junctions[3]. Details on sample fabrication and circuit parameters are found in Supplementary Note A. The sample was cooled to a temperature of 25 - 30 mK and measured in a standard setup as detailed in Supplementary Note B.
The qubit showed energy relaxation times $T_{1}$ between 5 to 8 µs at operation frequencies between 5 and 5.5 GHz. Similarly fabricated qubits without on-chip gate electrodes achieved only slightly longer $T_{1}$-times between 10 and 20 µs [3, 31], and these were mostly limited by TLS near circuit electrodes [30, 15]. However, the observed $T_{1}$-time falls within the estimate range of the radiative loss via the capacitive coupling to the four electrodes, which were placed in close vicinity to the qubit island to enhance the spatial resolution in TLS localization. This loss channel can be mitigated with an improved design of the qubit and on-chip electrodes [5]. Further details on loss are discussed in Supplementary Note C.
TLS spectroscopy
Figure 2: Finding the location of a TLS. a TLS spectroscopy using the protocol in the inset, to reveal the resonance frequencies of TLS from minima (dark pixels) in the resulting qubit population $P(|1\rangle)$. In each segment, the voltage on a different electrode $\alpha..\delta$ is increased by 1V. The highlighted trace shows a TLS as it is tuned through its symmetry point (arrow) according to Eq. (1) as illustrated in the lower inset. The TLS’ response strengths $\gamma_{i}$ to the different electrodes are obtained by fitting such traces to Eq. (1), and provide information about the TLS’ distance to the electrodes. b Difference between the measured TLS response strength ratio $\gamma_{i}/\gamma_{j}$ and the corresponding simulated E-field ratio (colorscale) for the TLS observed in a. Minima (dark pixels) indicate possible TLS positions. Each panel shows data from a different electrode pair as marked in the legends together with the measured tuning ratio. c The colorscale shows the difference sum $\sigma$ (Eq. 3) over all 6 unique combinations of electrode pairs. The minimum (white circle) marks the most probable TLS position.
@@
To measure the TLS’ coupling strengths to the four DC-electrodes, their resonances are traced by TLS spectroscopy while the voltages on the DC-electrodes are swept. As an example, Figure 2a shows the resonance of a TLS that was tuned through the symmetry point of its hyperbola Eq. (1), where in each segment the voltage on the indicated gate electrode $\alpha...\delta$ was stepwise increased by 1 V while other voltages were kept constant. The slope of the hyperbola in the different segments then depends on the tuning strength of the TLS by the corresponding electrode. It is obtained by fitting such traces to Eq. (1), where the factor for the induced asymmetry energy $2\textbf{p}\cdot\textbf{E}$ is replaced by $\sum_{i}\gamma_{i}V_{i}$. Here, $i\in\{\alpha,\beta,\gamma,\delta\}$ indicates the electrode that is biased by the voltage $V_{i}$, and $\gamma_{i}$ are the fitting tuning strengths that contain information on the distance of the TLS to the corresponding electrodes.
TLS localization
@@
To be able to detect a TLS in qubit $T_{1}$-swap-spectroscopy, its (resonant) coupling strength to the qubit $g=\left(\frac{\Delta}{\hbar\omega_{\mathrm{TLS}}}\right)\ \mathbf{p}\cdot\mathbf{E_{\mathrm{rms}}}$ must be large enough to result in a measurable decrease of the qubit’s energy relaxation rate. Thus, TLS can only be detected in circuit regions where the qubit’s AC-electric field strength $|E_{\mathrm{rms}}|$ exceeds $g_{\mathrm{min}}/p_{\parallel}$. Assuming $\Delta/\hbar\omega_{\mathrm{TLS}}\approx 1$ (most strongly coupling TLS near their symmetry point) and a field-parallel dipole moment of $p_{\parallel}\approx 1e$Å [34, 1, 9, 27], the required minimum coupling strength $g_{\mathrm{min}}$ can be estimated from the energy relaxation rate of the resonantly coupled qubit-TLS system [1] $\Gamma_{1}=2\,(g_{\mathrm{min}}/\hbar)^{2}/\Gamma+\Gamma_{1,Q}$, where $\Gamma_{1,Q}$ is the energy relaxation rate of the isolated qubit, and $\Gamma=\Gamma_{1,\mathrm{TLS}}/2+\Gamma_{2,\mathrm{TLS}}+\Gamma_{1,\mathrm{Q}}/2+\Gamma_{2,\mathrm{Q}}$ is the sum of TLS and qubit energy relaxation and dephasing rates. The assumption that TLS are detected if they reduce the qubit’s $T_{1}$ time by a factor of $\kappa$ translates into a minimum coupling strength of $g_{\mathrm{min}}=\hbar\sqrt{\kappa\cdot\Gamma_{\mathrm{1,Q}}\,\Gamma/2}$ and corresponding minimum AC-electric qubit field strength $E_{\mathrm{min}}=g_{\mathrm{min}}/p_{\parallel}$. A plot of these relations can be found in Supplementary Note E. For a qubit $T_{1}=7\$µs, TLS coherence times in the range of $T_{\mathrm{1,TLS}}\approx T_{\mathrm{2,TLS}}\approx 0.1-2\,$µs [1, 49, 30, 32, 11, 54], and $\kappa=5\%$, we find a $E_{\mathrm{min}}$ range of $\approx 0.3-3$ V/m.
@@
Figure 3a shows the simulated magnitude of the qubit’s AC-electric field $|E_{\mathrm{rms}}|$. A cross section through the edges of ground plane and qubit island is shown in Fig 3b and illustrates that in our qubit sample, TLS can only be detected within an $\approx 1-2$ µm distance from the edge of qubit electrodes where $|E_{\mathrm{rms}}|>E_{\mathrm{min}}$. Since $E_{\mathrm{min}}$ decreases with the square root of the qubit’s $T_{1}$ time, more coherent qubits are affected by TLS in a wider area which includes more weakly coupled ones as illustrated in Supplementary Note E.
@@
To show that the electric fields from different on-chip electrodes are indeed sufficiently parallel, such that the dipole moment in Eq. (2) can be canceled in order to simplify the solution for possible TLS positions, in Fig. 3c we plot the electric fields’ $\vec{x}$- and $\vec{z}$-components along the white line in Fig. 3a. Near the center of the gap between ground and qubit island, the E-fields point in different directions. However, within the short distance to the electrode edges where TLS are detected in our experiment, all E-fields are well aligned since their vector products with the qubit field $\vec{E_{q}}$ approach unity as plotted in Fig. 3d. This is also the case near the junction leads as shown in Supplementary Note E.
@@
Figure 4: TLS locations and properties. a Map of the individual positions of detected surface-TLS (yellow circles). Most TLS ($\approx$ 59%) were found at the Josephson junctions leads (see lower inset for a zoom on the DC-SQUID). Near the edges of capacitor island and ground plane, 27% and 14% of TLS were identified, respectively. b Histograms and cumulative distributions (insets) of the TLS’ electric dipole moment component $p_{\parallel}$ (upper panel) and TLS-qubit coupling strengths $g/h$ (lower panel) as estimated from the TLS’ positions, their coupling strengths to the electrodes, and the local magnitude of the qubit AC field $E_{\mathrm{rms}}$. c Percentage of TLS identified on the SQUID vs. ground plane plus island, plotted as a function of the electric field threshold $E_{\mathrm{min}}$ that accounts for TLS observability in swap spectroscopy. The solid lines are the energy participation ratios (EPRs) of the SQUID (red) and the ground plus island (blue) in regions where $E_{\mathrm{rms}}>E_{\mathrm{min}}$. The dashed lines are a fit of these ratios to the observed TLS distribution, obtained by assuming that the TLS density in the SQUID area is enhanced by a factor of 2. d Determined median TLS electric dipole moment $p_{\parallel}$ as a function of $E_{\mathrm{min}}$.
@@ Results
TLS distribution
To generate a map of the individual TLS positions on the qubit circuit, we swept the voltages on each of the four electrodes in a range between -60V and +100V and performed TLS spectroscopy in a 150-MHz wide window as shown in Fig. 2a. In total, 55 TLS were observed in a single qubit sample whose tuning strengths to all four electrodes could be characterized (see Supplementary Note F for further details). For the analysis, TLS were discarded whose coupling strength to the upper electrodes $\alpha$ and $\gamma$ were indistinguishable from zero so that their location could not be precisely determined. This is the case for TLS that reside on the long arm of the qubit island (see Fig. 1b).
@@
To check the mapping procedure’s reliability, the analysis is repeated for various values of $E_{\mathrm{min}}$. The resulting percentages of TLS found on the SQUID leads vs. those on the qubit island and ground plane are plotted in Fig. 4c. These remain at a ratio of 60:40 within the most probable range of $E_{\mathrm{min}}\approx 0.3..3$ V/m as estimated above, which backs the robustness of the method. Above a value of $E_{\mathrm{min}}\approx 6$ V/m, the procedure fails since it is forced to place all TLS on the DC-SQUID. For the given data and qubit sample, we estimate an uncertainty in the TLS positions of about 6 µm as further discussed in Supplementary Note G. This resolution can be enhanced with more suitable qubit designs that avoid sharp corners.
Excess TLS density near the SQUID
Notably, the observed TLS distribution points towards an excess density of TLS near the DC-SQUID leads. Since the probability to detect a surface-TLS scales with the square of the local qubit field, the expected ratio of the numbers of TLS observed near the SQUID vs. ground plane and qubit island can be estimated by comparing their energy participation ratios (EPRs). By calculating the EPR, we limit the integrals over $|E_{\mathrm{rms}}|^{2}$ to regions where $E_{\mathrm{rms}}>E_{\mathrm{min}}$ to account for TLS detectability in swap spectroscopy, which sets this calculation apart from the common participation ratio analysis [52] that includes also weakly coupled TLS. Further details are given in Supplementary Note E. The result is plotted with solid red and blue lines in Fig. 4c, and predicts that most TLS (>50%) would be observed on either the qubit island or ground plane in the whole range of reasonable values $E_{\mathrm{min}}<4$ V/m, in stark contrast to the experimental data.
@@ Discussion
We have demonstrated a method to determine the individual positions and electric dipole moments of TLS defects in a transmon qubit. The majority ($\approx 58\%$) of observed strongly interfering TLS in the qubit sample were found to reside on the qubit’s DC-SQUID. This confirms that the leads of tunnel junctions, due to their large electric energy participation, are a critical component which can dominate qubit loss [18, 50, 54]. This advocates for the use of wire-tapering techniques to dilute the electric field [35], and to minimize the size of lift-off structures.
Our results additionally indicate that the TLS density near the junction leads is enhanced by a factor of $\approx 2$. This may be attributed to their different fabrication technique as compared to the qubit capacitor, which can promote TLS formation in various ways. For example, junction formation by shadow evaporation and electron-beam lithography is associated with larger amounts of resist residuals and enhanced roughness of junction lead interfaces [20, 38, 24]. The junctions were fabricated with a lift-off process which reportedly leaves excess residues [45, 50]. Also, thinner films showed a larger density of grain boundaries associated with enhanced oxygen diffusion [7].
@@
Control over the local DC-electric field also enables one to actively suppress decoherence by tuning dominating TLS defects out of the qubit resonance [31, 10, 14]. For this, multiple electrodes provide independent control over TLS at different locations which enhances the ability to decouple the qubit from the decohering bath.
The current method cannot determine at which interfaces TLS are residing. This is possible with additional gate electrode placed above and below the qubit chip, allowing one to also distinguish TLS located at the metal-substrate interface from those at the metal-air interface and the substrate as we showed in previous work [4]. Moreover, the combination with TLS tuning by applied mechanical strain provides additional information of TLS densities in tunnel junction barriers [30], and is useful to enhance the numbers of observable TLS in a given qubit circuit.
@@
Data availability
The raw data that support the findings of this study are available from the public repository Zenodo, DOI:10.5281/zenodo.18847452.
@@ Acknowledgements
The authors declare no competing financial or non-financial interests.
@@
[…]
Chip dicing. The wafers are finally diced into chips of 6x6 mm size after additional coating with protective resist as described above.
Figure S1 shows a photograph of a finished chip. A summary of the qubit circuit parameters is given by Table S1.
Fig. S1. Photograph of the qubit chip for TLS mapping.
@@
Table S1. Measured and simulated fabrication parameters of the studied qubit sample. Capacitance values were extracted with Ansys Maxwell. Junction (JJ) and SQUID resistances are averages from 4-point measurements at room temperature.
@@
$Q_{l}=\left(\frac{C_{\mathrm{tot}}}{C_{c}}\right)^{2}\frac{Z_{q}}{\mathrm{Re}(Z_{\mathrm{eff}})},$ (A1)
where $C_{\mathrm{tot}}=C_{q}+\sum_{i}C_{q,i}$ is the total capacitance, and $Z_{q}=\sqrt{L_{q}/C_{\mathrm{tot}}}\approx 290\,\Omega$ is the qubit impedance estimated from the inductance of its two parallel Josephson junctions $L_{q}=0.5\cdot L_{J}=0.5\cdot h/4\pi eI_{c}$ each having a critical current of $I_{c}\approx 23$ nA as shown in Table Supplementary Note A. Assuming a real part impedance of $\mathrm{Re}(Z_{\mathrm{eff}})\approx 50\,\Omega$ for the coplanar gate electrode including its feed line results in $Q_{l}\approx 0.6$ million. The resulting loaded quality factor for all four electrodes is $Q_{\mathrm{tot}}=(4/Q_{l})^{-1}$ and limits the qubit’s energy relaxation time to $T_{1}=Q_{\mathrm{tot}}/2\pi f_{q}\approx 5\,$µs at a qubit resonance frequency of $f_{q}=5$ GHz.
@@
Supplementary Note D Electric field simulations
For the TLS mapping procedure and the energy-participation-ratio (EPR) analysis, the AC-electric field distribution of the transmon qubit was simulated using ANSYS HFSS’s Eigenmode Solver. The DC-electric fields emitted by the four gate electrodes were simulated with ANSYS MAXWELL, separately biasing each electrode with 1V. In all models, the simulation volume was restricted to a $\times1mm$$\times$1mm large region centered around the qubit. All thin-film structures, including the qubit island, Josephson junctions, gate electrodes, resonator, bias line, and ground plane, were modeled as 2D sheets on a $\times1mm$$\times$0.5mm sapphire substrate. Except for the junctions, all sheets were assigned the “Perfect E” boundary condition, while the junction sheets were assigned the “Lump RLC” boundary condition. Assuming a relative permittivity $\epsilon_{\mathrm{r}}=10$ and thickness d = 2nm of the junction barrier, the applied junction capacitances $C_{\mathrm{JJ}}$ were estimated from the known junction area A, using $C_{\mathrm{JJ}}=A\epsilon_{0}\epsilon_{\mathrm{r}}/d$, to approximately 3.6 fF. The junction inductances were adjusted such that the simulated eigenfrequency of the transmon qubit was approximately $f_{\mathrm{q}}\approx$ 5 GHz. Along the edges of qubit sheets and the entire SQUID loop, the precision of the adaptive mesh solver was set to a minimum of 100 nm. Finally, the energy of the qubit mode was normalized to one photon to obtain the electric field strength $E_{\mathrm{rms}}$.
Systematic offsets in the simulated electric field strength, e.g. due to averaging over simulation mesh elements, play a minor role since only relative field strengths are considered in the method. We used the same mesh for all simulations and increased the meshing resolution in a 2 µm wide region around electrode edges. Systematic errors however do affect the estimation of the TLS dipole moments, which are directly calculated from the simulated local DC-field at the deduced TLS position and the measured TLS tuning strength. The steep gradient of $E_{\mathrm{rms}}$ near electrode edges enhances the uncertainty of the TLS dipole moment. Despite this, the obtained TLS dipole moments (Fig. 4b in the main manuscript) have a reasonable mean value and do not show excessive spread compared to other results [27].
Supplementary Note E TLS detectability limits
@@
$\Gamma_{1}=\frac{2\,(g/h)^{2}\,\Gamma}{(\Gamma/2\pi)^{2}+\delta^{2}}+\Gamma_{\mathrm{1,q}}.$ (A2)
@@
Solving for the coupling strength $g$ where Eq. (A2) predicts that the qubit’s energy relaxation time at resonance $\delta=0$ is decreased by a factor $\kappa$ then results in the minimum coupling strength $g_{\mathrm{min}}=\hbar\sqrt{\kappa\cdot\Gamma_{\mathrm{1,q}}\,\Gamma/2}$ and corresponding minimum AC-electric qubit field strength $E_{\mathrm{min}}=g_{\mathrm{min}}/p_{\parallel}$.
@@
Energy participation ratio (EPR). The probability to detect a TLS increases with the square of the local qubit AC-field, since the qubit decay rate $\Gamma_{1}$, which is used as the signal in swap spectroscopy, scales with the qubit-TLS coupling strength $g^{2}\propto|E_{\mathrm{rms}}|^{2}$ (see Eq. A2 and definition of $g$ in the main text). This is analogous to surface loss analyses which consider the energy participation ratio (EPR) of different circuit areas, i.e. the fraction of the qubit energy that is stored there [53]. We thus estimate the expected ratio of the numbers of TLS observed on the SQUID vs. those on the qubit capacitor by comparing integrals over $|E_{\mathrm{rms}}|^{2}$ in the vicinity of these elements.
For this, the simulated $|E_{\mathrm{rms}}|$ is exported on a rectangular grid, and the sum of $|E_{\mathrm{rms}}|^{2}$ is calculated for pixels inside the SQUID and capacitor areas whose dimensions are shown in Fig. S5. Moreover, this sum is limited to regions where $|E_{\mathrm{rms}}|>E_{\mathrm{min}}$, in order to account for the detectability of TLS in swap spectroscopy. The resulting EPRs of SQUID and ground+island are shown in the main text’s Fig. 4c as a function of the chosen value of $E_{\mathrm{min}}$.
Fig. S5: Areas assigned to the ground plane (blue), the qubit island (red), and the SQUID (indicated square), as used for calculating energy participation ratios. The white lines indicate electrode edges.
Field directions. The method to find TLS positions requires that the gate electrode fields are parallel at the position of the TLS. This is confirmed with Fig. S7 also for the region near the Josephson junctions. Fig. S7a shows the strength of the qubit’s AC-electric field $|E_{\mathrm{rms}}|$ and a zoom onto the DC-SQUID. A vertical cross-section is shown in Fig. S7b, where the junction leads are indicated by gray bars that are separated by 6 µm. The region where $|E_{\mathrm{rms}}|>E_{\mathrm{min}}$ (here for $E_{\mathrm{min}}$=0.75 V/m) is shaded red. Directly at the junctions, the detection region extends up to 1.8 µm distance from the lead edges. Similar to Fig. 3 in the main text, the $\vec{x}$- and $\vec{z}$-directions of the electric fields are plotted in Fig. S7c, and their vector products shown in Fig. S7d indicate that they are practically parallel within the vicinity of junction leads.
Fig. S7: Strength and orientations of AC- and DC-electric fields near the DC-SQUID. a Magnitude of the qubit’s AC-electric field $|E_{\mathrm{rms}}|$ as simulated with Ansys HFSS, with a zoom onto the DC-SQUID. b Cross-section of $|E_{\mathrm{rms}}|$ along the black line in a. TLS can be detected in the red-shaded area where the field exceeds a minimum strength $E_{\mathrm{min}}$. c Components of the electric fields from the four electrodes and the qubit, plotted in $\vec{x}$- and $\vec{z}$ directions (left and right panel). d The normalized vector product between the fields of gate electrodes and the qubit indicates that they are practically parallel in the range where TLS can be detected.
Supplementary Note F Data acquisition and TLS properties
@@
Supplementary Note G Resolution in TLS mapping